# Characterizing higher Auslander(-Gorenstein) Algebras

### v1

@@ 2. Higher Auslander-Gorenstein Algebras
The idea is to show that ${\mathrm{Gproj}}^{\leq n}(\Gamma)$ is the torsion-free class of some hereditary torsion pair, and so is closed under taking injective envelopes [Di66, Theorem 2.9].
Obviously the subcategory ${\mathrm{Gproj}}^{\leq n}(\Gamma)$ is closed under extensions. By Lemma 2.3 ($b^{\prime}$), it is also closed under taking submodules. Hence, the subcategory ${\mathrm{Gproj}}^{\leq n}(\Gamma)$ is closed under taking extensions and submodules and so is the torsion-free class of some torsion pair $(\mathcal{T},{\mathrm{Gproj}}^{\leq n}(\Gamma))$ in ${\rm{{mod}}}(\Gamma)$ [Di66, Theorem 2.3].
We now show that ${{}^{\perp_{0}}\Gamma}={{}^{{\perp}_{0}}{\mathrm{Gproj}}^{\leq n}(\Gamma)}=\mathcal{T}$. Let $X\in\mathcal{T}$. Then ${\rm{Hom}}_{\Gamma}(X,M)=0$ for $M=\Gamma\in{\mathrm{Gproj}}^{\leq n}(\Gamma)$, we get $X\in{{}^{\perp_{0}}\Gamma}$. This shows that ${{}^{\perp_{0}}\Gamma}\supseteq\mathcal{T}$.
On the other hand, suppose that $X\in{{}^{\perp_{0}}}\Gamma$. By ${}^{\perp_{0}}\Gamma\subseteq{}^{\perp_{n}}\Gamma$, ${\rm{Ext}}^{i}_{\Gamma}(X,\Gamma)=0$ for every $0\leq i\leq n$. Hence, if $Y$ is a projective $\Gamma$-module, then ${\rm{Ext}}^{i}_{\Gamma}(X,Y)=0$ for every $0\leq i\leq n$. Thus, by a dimension shifting argument, we deduce that ${\rm{Hom}}_{\Gamma}(X,Y)=0={\rm{Ext}}^{1}_{\Gamma}(X,Y)$ for each $Y\in{\mathrm{Proj}}^{\leq n-1}(\Gamma)$. We also deduce that ${\rm{Hom}}_{\Gamma}(X,Y)=0$ for each $Y\in{\mathrm{Proj}}^{\leq n}(\Gamma)$.
Also, if $Y$ is a Gorenstein projective $\Gamma$-module, then $Y$ is a submodule of a projective $\Gamma$-module and so ${\rm{Hom}}_{\Gamma}(X,Y)=0$.
Let now $Y$ be a $\Gamma$-module with ${\rm{Gpd}}({}_{\Gamma}Y)\leq n$. Then, by [Ho04a, Theorem 2.10], there exists an exact sequence
@@
such that $G$ is a Gorenstein projective module and $K$ is a $\Gamma$-module of projective dimension less than $n$. By applying ${\rm{Hom}}_{\Gamma}(X,-)$ to this exact sequence, we deduce that ${\rm{Hom}}_{\Gamma}(X,Y)=0$ for each $Y\in{\mathrm{Gproj}}^{\leq n}(\Gamma)$. We have shown that ${{}^{\perp_{0}}\Gamma}=\mathcal{T}$.
@@
This finishes the proof of Step 2.
@@
By Step II, for every $X\in{\mathrm{Gproj}}^{\leq n}(\Gamma)$, $\mathbf{E}(X)\in{\mathrm{Gproj}}^{\leq n}(\Gamma)$. On the other hand, as mentioned in Subsection 1.11, by [Ho04b, Theorem 2.2], the projective dimension of $\mathbf{E}(X)$ equals to its Gorenstein projective dimension and so $\mathbf{E}(X)\in\mathrm{proj}^{\leq n}(\Gamma)$.
@@
Propositions 2.1 and 2.4 could be enhanced to infinitely generated modules, that is, we could show the following result. Since its proof is analogous to that of Propositions 2.1 and 2.4, we only give a sketch of proof.
@@
$\Rightarrow$”. Let $\Gamma$ be an $n$-minimal Auslander-Gorenstein algebra. Let $X$ be an arbitrary injective $\Gamma$-module. By Facts 1.11 $(b)$, $X$ is a direct sum of indecomposable injective $\Gamma$-modules, the latter being finitely generated over an Artin algebra. From this fact, we deduce that each injective $\Gamma$-module is a direct sum of finitely generated injective $\Gamma$-modules.
@@
$\Leftarrow$”. For this direction, the proof is nearly the same as the proof of the implication $\Leftarrow$” in Proposition 2.4 except that in Step II, one needs to observe that the subcategory ${\mathrm{GProj}}^{\leq n}(\Gamma)$ is also closed under taking direct products and to use [Di66, Theorem 2.3] to conclude. ∎
Now by combining Propositions 2.1, 2.4, and 2.5 we have the following theorem that can be considered as a higher generalization of Auslander-Tachikawa theorem.
@@
Based on the proof of Theorem 2.6 (in fact, Propositions 2.1, 2.4, and 2.5) we have the following results.
@@
In Step 2 of the proof of Proposition 2.4, we showed that the pair $({{}^{\perp_{0}}\Gamma},{\mathrm{Gproj}}^{\leq n}(\Gamma))$ is a hereditary torsion pair. A similar argument shows that the pair $({{}^{\underline{\perp}_{0}}\Gamma},{\mathrm{GProj}}^{\leq n}(\Gamma))$ is also a hereditary torsion pair. ∎
@@
Kong used of the notion of submodule categories and characterized Artin algebras whose Gorenstein projective modules form an abelian category; see [Ko14, Corollary 2.3]. In the following, we will prove a similar property for higher Auslander-Gorenstein algebras.
@@
A combination of our results and some classical results about torsion pairs gives us a better perspective about higher Auslander-Gorenstein algebras, see Corollaries 2.10 and 2.11.
@@
By Corollary 2.7, the pair $({{}^{\perp_{0}}\Gamma},{\mathrm{Gproj}}^{\leq n}(\Gamma))$ is a torsion pair. By [AR94, Proposition 5.8], ${\mathrm{Gproj}}^{\leq n}(\Gamma)$ is a functorially finite subcategory of ${\rm{{mod}}}(\Gamma)$ and so by Theorem 1.4 that we recalled it in the Subsection 1.5, all the statements hold and also are equivalent with each other. Note that $\Gamma\in{\mathrm{Gproj}}^{\leq n}(\Gamma)$ and so $\mathrm{Ann}_{\Gamma}({\mathrm{Gproj}}^{\leq n}(\Gamma))=0$. But, by [Sm84, Lemma 0.2], $\mathrm{Ann}_{\Gamma}({{}^{\perp_{0}}\Gamma})\neq 0$. ∎
@@
By Corollary 2.7, the pair $({{}^{\perp_{0}}\Gamma},{\mathrm{Gproj}}^{\leq n}(\Gamma))$ is a torsion pair and so, by [Ho82, Lemmas 1,2, and 3], all the statements hold; see also [AS81, Corollaries 3.4 and 3.7]. ∎
@@
By [Ho82, Lemmas 2 and 3], all the statements hold; see also [AS81, Corollaries 3.4 and 3.7]. ∎
3. Selfinjective and $\tau_{n}$-selfinjective Algebras
As an application of our results in Section 2, in this section we study selfinjective and $\tau_{n}$-selfinjective algebras and characterize them in a categorical sense.
We first have the following result that gives a categorical description for when an $n$-minimal Auslander-Gorenstein algebra $\Gamma$ is self-injective. Recall that an Artin algebra $\Gamma$ is said to be selfinjective if it is injective as a left $\Gamma$-module.
Proposition 3.1.
Let $n$ be a positive integer and an Artin algebra $\Gamma$ be an $n$-minimal Auslander-Gorenstein algebra. Then the following statements are equivalent:
  • $(a)$
    $\Gamma$ is selfinjective;
  • $(b)$
    ${\mathrm{Gproj}}^{\leq n}(\Gamma)$ is an abelian category;
  • $(c)$
    ${{}^{\perp_{0}}\Gamma}\subseteq{{}^{\perp_{n+1}}\Gamma}$;
  • $(d)$
    ${\mathrm{GProj}}^{\leq n}(\Gamma)$ is an abelian category;
  • $(e)$
    ${{}^{\underline{\perp}_{0}}\Gamma}\subseteq{{}^{\underline{\perp}_{n+1}}\Gamma}$.
Proof.
$(a)\Leftrightarrow(b)$. Let ${\mathrm{Gproj}}^{\leq n}(\Gamma)$ be an abelian category and $\Gamma$ be non-selfinjective. Then ${\rm{Ggldim}}(\Gamma)$ $=n+1$, so there exists a finitely generated $\Gamma$-module $X$ such that ${\rm{Gpd}}({}_{\Gamma}X)=n+1$. Let
$0\to K\stackrel{{\scriptstyle\kappa}}{{\to}}P_{0}\stackrel{{\scriptstyle\varepsilon}}{{\longrightarrow}}X\longrightarrow 0$
be a short exact sequence with $P_{0}\in\mathrm{proj}(\Gamma)$. Then $K\in{\mathrm{Gproj}}^{\leq n}(\Gamma)$. If $\pi:P_{0}\longrightarrow Y$ is the cokernel of the natural monomorphism $\kappa$ in ${\mathrm{Gproj}}^{\leq n}(\Gamma)$, then the sequence
$0\longrightarrow K\stackrel{{\scriptstyle\kappa}}{{\longrightarrow}}P_{0}\stackrel{{\scriptstyle\pi}}{{\longrightarrow}}Y$
is an exact sequence in ${\mathrm{Gproj}}^{\leq n}(\Gamma)$ as ${\mathrm{Gproj}}^{\leq n}(\Gamma)$ is an abelian category. Since $\Gamma\in{\mathrm{Gproj}}^{\leq n}(\Gamma)$, the sequence
$0\longrightarrow{\rm{Hom}}_{\Gamma}(\Gamma,K)\longrightarrow{\rm{Hom}}_{\Gamma}(\Gamma,P_{0})\longrightarrow{\rm{Hom}}_{\Gamma}(\Gamma,Y)$
of abelian groups is also exact and so the sequence
$0\longrightarrow K\stackrel{{\scriptstyle\kappa}}{{\longrightarrow}}P_{0}\stackrel{{\scriptstyle\pi}}{{\longrightarrow}}Y$
is also exact in ${\rm{{mod}}}(\Gamma)$. Thus the induce map $\overline{\pi}:X\simeq P_{0}/K\longrightarrow Y$ is a monomorphism. Hence, $X$ is a submodule of $Y$ and so there exists a short exact sequence
$0\longrightarrow X\longrightarrow Y\longrightarrow Z\longrightarrow 0$
in ${\rm{{mod}}}(\Gamma)$ such that ${\rm{Gpd}}({}_{\Gamma}Y)\leq n$. Since ${\rm{Ggldim}}(\Gamma)=n+1$, ${\rm{Gpd}}({}_{\Gamma}Z)\leq n+1$ and so ${\rm{Gpd}}({}_{\Gamma}X)\leq n$, which is a contradiction.
We have shown that if ${\mathrm{Gproj}}^{\leq n}(\Gamma)$ is an abelian category, then $\Gamma$ is selfinjective. The other direction is trivial.
$(a)\Leftrightarrow(c)$. Assume that ${{}^{\perp_{0}}\Gamma}\subseteq{{}^{\perp_{n+1}}\Gamma}$. Suppose that $\Gamma$ has Gorenstein global dimension $n+1$. We first show that ${{}^{\perp_{n+1}}\Gamma}\subseteq{\mathrm{Gproj}}^{\leq n}(\Gamma)$. Let $X\in{}^{\perp_{n+1}}\Gamma$. Then ${\rm{Ext}}_{\Gamma}^{n+1}(X,\Gamma)=0$. Since ${\rm{Ggldim}}(\Gamma)=n+1$, ${\rm{Gpd}}({}_{\Gamma}X)\leq n$ and hence ${}^{\perp_{n}}\Gamma\subseteq{\mathrm{Gproj}}^{\leq n}(\Gamma)$. Then ${{}^{\perp_{0}}\Gamma}\subseteq{}^{\perp_{n+1}}\Gamma\subseteq{\mathrm{Gproj}}^{\leq n}(\Gamma)$. By Corollary 2.7, $({{}^{\perp_{0}}\Gamma},{\mathrm{Gproj}}^{\leq n}(\Gamma))$ is a torsion pair in ${\rm{{mod}}}(\Gamma)$ and so ${{}^{\perp_{0}}\Gamma}=0$. Therefore, ${\mathrm{Gproj}}^{\leq n}(\Gamma)={\rm{{mod}}}(\Gamma)$, which contradicts the assumption ${\rm{Ggldim}}(\Gamma)=n+1$. So ${{}^{\perp_{0}}\Gamma}\subseteq{{}^{\perp_{n+1}}\Gamma}$ implies that $\Gamma$ is selfinjective.
The other direction is trivial.
$(a)\Leftrightarrow(d)$. The proof is similar to that of $(a)\Leftrightarrow(b)$.
$(a)\Leftrightarrow(e)$. The proof is similar to that of $(a)\Leftrightarrow(c)$. ∎
Now we can give a categorical description for the selfinjectivity of an Artin algebra.
Corollary 3.2.
Let $\Gamma$ be an Artin algebra. Then $\Gamma$ is selfinjective if and only if the categories ${\mathrm{Gproj}}(\Gamma)$ and ${\mathrm{Gproj}}^{\leq 1}(\Gamma)$ are abelian categories if and only if the categories ${\mathrm{GProj}}(\Gamma)$ and ${\mathrm{GProj}}^{\leq 1}$ $(\Gamma)$ are abelian categories.
Proof.
Let the categories ${\mathrm{Gproj}}(\Gamma)$ and ${\mathrm{Gproj}}^{\leq 1}(\Gamma)$ be abelian categories. By [Ko14, Theorem 2.1], $\Gamma$ is an Auslander algebra and hence by Proposition 3.1, $\Gamma$ is selfinjective. For the second equivalence, the proof is similar. ∎
We now turn our attention to $\tau_{n}$-selfinjective algebras. Let us recall their definition.
@@ 3. Selfinjective and $\tau_{n}$-selfinjective Algebras
Let $\Lambda$ be an Artin algebra and $\mathcal{X}$ be a full subcategory of ${\rm{{mod}}}(\Lambda)$. The subcategory $\mathcal{X}$ is cogenerating if for every object $M\in{\rm{{mod}}}(\Lambda)$ there exist an object $X\in\mathcal{X}$ and a monomorphism $M\longrightarrow X$. The concept of generating subcategories is defined dually. We also call $\mathcal{X}$ $n$-rigid if ${\rm{Ext}}^{i}_{\Lambda}(\mathcal{X},\mathcal{X})=0$ for each $0<i<n$.
Let now $\tau_{n}$ and $\tau_{n}^{-}$ be the $n$-Auslander-Reiten translations; see [Iy07a, Subsection 1.4] or [Iy11, Definition 1.1]. A generating-cogenerating $n$-rigid subcategory $\mathcal{X}$ of ${\rm{{mod}}}(\Lambda)$ is called $n$-precluster tilting subcategory if it is functorially finite, $\tau_{n}(\mathcal{X})\subseteq\mathcal{X}$, and $\tau_{n}^{-}(\mathcal{X})\subseteq\mathcal{X}$ [IS18, Definition 3.2]. Moreover, if $\mathcal{X}$ admits an additive generator $M$, we say that $\mathcal{X}$ is a finite $n$-precluster tilting subcategory and $M$ is an $n$-precluster tilting module. By following Iyama-Solberg [IS18, Definition 3.4], Artin algebras which have $n$-precluster tilting modules are also called $\tau_{n}$-selfinjective.
@@
Now, by combining Iyama-Solberg correspondence and our results (Theorem 2.6), we can prove a result that gives us a new characterization of the class of Artin algebras having $n$-precluster tilting modules, i.e. $\tau_{n}$-selfinjective algebras. We need the following lemma which is of independent interest and follows from Lemmas 1.5 and 1.6.
@@
    The functor $\mathrm{Hom}_{\Lambda}(M,-):{\rm{{mod}}}(\Lambda)\to{\rm{{mod}}}(\Gamma)$ restricts to an equivalence $\mathrm{add}(M)\simeq\mathrm{proj}(\Gamma)$, which further restricts to another equivalence $\mathrm{inj}(\Lambda)\simeq\mathrm{proj}(\Gamma)\cap\mathrm{inj}(\Gamma)$, so $\mathrm{add}(\mathcal{Q})$ $=\mathrm{proj}(\Gamma)\cap\mathrm{inj}(\Gamma)$.
@@
The statement $(a)$ can be proved as in [ARS, chapter 5, Lemma 5.3]. For $(b)$, the fact that $M_{\Gamma}$ is projective follows from the duality $\mathrm{Hom}_{\Lambda}(-,M)$ between $\mathrm{add}(M)$ and $\mathrm{proj}(\Gamma^{op})$ and the fact that ${}_{\Lambda}\Lambda\in\mathrm{add}(M)$.
@@
For $(c)$, the two functors $\mathrm{Hom}_{\Gamma}(I,-)$ and $D\mathrm{Hom}_{\Gamma}(-,Q)$ from ${\rm{{mod}}}(\Gamma)$ to ${\rm{{mod}}}(\Lambda)$ are right exact (in fact exact), so by Watts’ theorem, one need to verify that they coincident with $M\otimes_{\Gamma}-$ on the left regular module ${}_{\Lambda}\Lambda$, which is obvious by $Q=\mathrm{Hom}_{\Lambda}(M,D\Lambda)\simeq D(M)$ and $I\simeq\mathrm{Hom}_{\Gamma}(M,\Gamma)\simeq\nu^{-1}Q$. The existence of the adjoint triple follows from the usual tensor-Hom adjunction.
@@
By the first lemma of [Au71, Chapter III, Section 4, Page 48]. this equivalence can be extended to an equivalence ${\rm{{mod}}}(\Lambda)\simeq\mathrm{sub}^{2}(\mathcal{Q})$, as $Q$ is injective. This proves $(d)$.
For $(d^{\prime})$, by Fact 1.11 $(a)(b)$, the equivalence $\mathrm{inj}(\Lambda)\simeq\mathrm{add}(Q)$ can be extended to ${\mathrm{Inj}}(\Lambda)\simeq\mathrm{Add}(Q)$. Again the latter can be extended to ${\rm{Mod}}(\Lambda)\simeq\mathrm{Sub}^{2}(\mathcal{Q})$. ∎
@@
$(b)\Rightarrow(a)$. By Proposition 2.4, $\Gamma$ is an $n$-minimal Auslander-Gorenstein algebra and so by Iyama-Solberg correspondence [IS18, Theorem 4.5], there is an Artin algebra $\Lambda^{\prime}$ with $n$-precluster tilting $\Lambda^{\prime}$-module $M^{\prime}$ such that ${\rm{{mod}}}(\Lambda^{\prime})\simeq{\mathrm{Gproj}}^{\leq n-1}(\Gamma)\simeq{\rm{{mod}}}(\Lambda)$. Hence, $\Lambda$ and $\Lambda^{\prime}$ are also Morita equivalent.
@@
Lemmas 1.6 and 3.4, allowed us to generalize our results from finitely generated modules to all modules, see Theorems 2.6 and 3.5. Of course, we are not limited to these lemmas and can follow other ways.
@@
By [KS98, Proposition 3.5], there is a (locally Noetherian) Grothendieck category $\mathcal{A}$ and a functor $F:\mathcal{A}\longrightarrow{\rm{Mod}}(\Gamma)$ such that $F$ induces an equivalence between ${\mathrm{Inj}}(\mathcal{A})$ and ${\rm{Add}}(\mathcal{Q})$. By the proof of [KS98, Proposition 3.5], the functor $F$ is a composition of the right adjoint of a quotient functor with an exact functor and so $F$ is left exact. Hence, $F$ induces an equivalence between $\mathcal{A}$ and $\mathrm{Sub}^{2}(\mathcal{Q})$. On the other hand, by a similar argument like the proof of Proposition 2.1, ${\mathrm{GProj}}^{\leq n-1}(\Gamma)=\mathrm{Sub}^{2}(\mathcal{Q})$ and so ${\mathrm{GProj}}^{\leq n-1}(\Gamma)$ is an abelian category.

### v2

@@ 3. Higher Auslander-Gorenstein Algebras
The idea is to show that ${\mathrm{Gproj}}^{\leq n}(\Gamma)$ is the torsion-free class of some hereditary torsion pair, and so is closed under taking injective envelopes [16, Theorem 2.9].
Obviously the subcategory ${\mathrm{Gproj}}^{\leq n}(\Gamma)$ is closed under extensions. By Lemma 3.3 ($b^{\prime}$), it is also closed under taking submodules. Hence, the subcategory ${\mathrm{Gproj}}^{\leq n}(\Gamma)$ is closed under taking extensions and submodules and so is the torsion-free class of some torsion pair $(\mathcal{T},{\mathrm{Gproj}}^{\leq n}(\Gamma))$ in ${\rm{{mod}}}(\Gamma)$ [16, Theorem 2.3].
We now show that ${{}^{\perp_{0}}\Gamma}={{}^{{\perp}_{0}}{\mathrm{Gproj}}^{\leq n}(\Gamma)}=\mathcal{T}$. Let $X\in\mathcal{T}$. Then ${\rm{Hom}}_{\Gamma}(X,\Gamma)=0$ as $\Gamma\in{\mathrm{Gproj}}^{\leq n}(\Gamma)$, and so $X\in{{}^{\perp_{0}}\Gamma}$. This shows ${{}^{\perp_{0}}\Gamma}\supseteq\mathcal{T}$.
On the other hand, suppose $X\in{{}^{\perp_{0}}}\Gamma$. By $^{\perp_{0}}\Gamma\subseteq{}^{\perp_{n}}\Gamma$, ${\rm{Ext}}^{i}_{\Gamma}(X,\Gamma)=0$ for every $0\leq i\leq n$. Hence, if $Y$ is a projective $\Gamma$-module, then ${\rm{Ext}}^{i}_{\Gamma}(X,Y)=0$ for every $0\leq i\leq n$. If $Y$ is a $\Gamma$-module with ${\rm{pd}}({}_{\Gamma}Y)=1$, then $Y$ has a projective resolution as the following.
$0\longrightarrow P_{1}\longrightarrow P_{0}\longrightarrow Y\longrightarrow 0$
Applying ${\rm{Hom}}_{\Gamma}(X,-)$ to this exact sequence gives the long exact sequence
[…]
and so ${\rm{Ext}}^{i}_{\Gamma}(X,Y)=0$ for every $0\leq i\leq n-1$. By applying this argument for finite steps, we deduce that ${\rm{Hom}}_{\Gamma}(X,Y)=0={\rm{Ext}}^{1}_{\Gamma}(X,Y)$ for each $Y\in\mathrm{proj}^{\leq n-1}(\Gamma)$. We also deduce that ${\rm{Hom}}_{\Gamma}(X,Y)=0$ for each $Y\in\mathrm{proj}^{\leq n}(\Gamma)$.
Let now $Y$ be a $\Gamma$-module with ${\rm{Gpd}}({}_{\Gamma}Y)\leq n$. Then, by [20, Theorem 2.10], there exists an exact sequence
@@
such that $G$ is a Gorenstein projective module and $K$ is a $\Gamma$-module of projective dimension less than $n$. By applying ${\rm{Hom}}_{\Gamma}(X,-)$ to this exact sequence, we deduce that ${\rm{Hom}}_{\Gamma}(X,Y)=0$ for each $Y\in{\mathrm{Gproj}}^{\leq n}(\Gamma)$ as $G$ is a submodule of a projective $\Gamma$-module and the projective dimension of $K$ is less than $n$. We have shown that ${{}^{\perp_{0}}\Gamma}=\mathcal{T}$.
@@
This finishes the proof of Step II.
@@
By Step II, for every $X\in{\mathrm{Gproj}}^{\leq n}(\Gamma)$, $\mathbf{E}(X)\in{\mathrm{Gproj}}^{\leq n}(\Gamma)$. On the other hand, by [21, Theorem 2.2], the projective dimension of $\mathbf{E}(X)$ equals to its Gorenstein projective dimension and so $\mathbf{E}(X)\in\mathrm{proj}^{\leq n}(\Gamma)$.
@@
Propositions 3.1 and 3.4 could be enhanced to infinitely generated modules, that is, we could show the following result. Since its proof is analogous to that of Propositions 3.1 and 3.4, we only give a sketch of proof.
@@
$\Rightarrow$”. Let $\Gamma$ be an $n$-minimal Auslander-Gorenstein algebra. Let $X$ be an arbitrary injective $\Gamma$-module. By [32, Theorem 3.2], $X$ is a direct sum of indecomposable injective $\Gamma$-modules, the latter being finitely generated over an Artin algebra. From this fact, we deduce that each injective $\Gamma$-module is a direct sum of finitely generated injective $\Gamma$-modules.
@@
$\Leftarrow$”. For this direction, the proof is nearly the same as the proof of the implication $\Leftarrow$” in Proposition 3.4 except that in Step II, one needs to observe that the subcategory ${\mathrm{GProj}}^{\leq n}(\Gamma)$ is also closed under taking direct products and to use [16, Theorem 2.3] to conclude. ∎
Now by combining Propositions 3.1, 3.4, and 3.5 we have the following theorem that can be considered as a higher generalization of Auslander-Tachikawa theorem.
@@
Based on the proof of Theorem 3.6 (in fact, Propositions 3.1, 3.4, and 3.5) we have the following results.
@@
In Step II of the proof of Proposition 3.4, we showed that the pair $({{}^{\perp_{0}}\Gamma},{\mathrm{Gproj}}^{\leq n}(\Gamma))$ is a hereditary torsion pair. A similar argument shows that the pair $({{}^{\underline{\perp}_{0}}\Gamma},{\mathrm{GProj}}^{\leq n}(\Gamma))$ is also a hereditary torsion pair. ∎
@@
Kong used of the notion of submodule categories and characterized Artin algebras whose Gorenstein projective modules form an abelian category; see [33, Corollary 2.3]. In the following, we will prove a similar property for higher Auslander-Gorenstein algebras.
@@
A combination of our results and some classical results about torsion pairs gives us a better perspective about higher Auslander-Gorenstein algebras, see Corollaries 3.10 and 3.11.
@@
By Corollary 3.7, the pair $({{}^{\perp_{0}}\Gamma},{\mathrm{Gproj}}^{\leq n}(\Gamma))$ is a torsion pair. By [4, Proposition 5.8], ${\mathrm{Gproj}}^{\leq n}(\Gamma)$ is a functorially finite subcategory of ${\rm{{mod}}}(\Gamma)$ and so by [41, Theorem] and [1, Proposition 1.1], all the statements hold and also are equivalent with each other. Note that $\Gamma\in{\mathrm{Gproj}}^{\leq n}(\Gamma)$ and so $\mathrm{Ann}_{\Gamma}({\mathrm{Gproj}}^{\leq n}(\Gamma))=0$. But, by [41, Lemma 0.2], $\mathrm{Ann}_{\Gamma}({{}^{\perp_{0}}\Gamma})\neq 0$. ∎
@@
By Corollary 3.7, the pair $({{}^{\perp_{0}}\Gamma},{\mathrm{Gproj}}^{\leq n}(\Gamma))$ is a torsion pair and so, by [22, Lemmas 1,2, and 3], all the statements hold; see also [6, Corollaries 3.4 and 3.7]. ∎
@@
By [22, Lemmas 2 and 3], all the statements hold; see also [6, Corollaries 3.4 and 3.7]. ∎
4. $\tau_{n}$-selfinjective Algebras
As an application of our results in Section 3, in this section we study $\tau_{n}$-selfinjective algebras and characterize them in a categorical sense. Let us recall their definition.
@@ 4. $\tau_{n}$-selfinjective Algebras
Let $\Lambda$ be an Artin algebra and $\mathcal{X}$ be a full subcategory of ${\rm{{mod}}}(\Lambda)$. The subcategory $\mathcal{X}$ is cogenerating if for every object $M\in{\rm{{mod}}}(\Lambda)$ there exist an object $X\in\mathcal{X}$ and a monomorphism $M\longrightarrow X$. The concept of generating subcategories is defined dually. Also $\mathcal{X}$ is called $n$-rigid if ${\rm{Ext}}^{i}_{\Lambda}(\mathcal{X},\mathcal{X})=0$ for each $0<i<n$.
Let now $\tau_{n}$ and $\tau_{n}^{-}$ be the $n$-Auslander-Reiten translations; see [26, Subsection 1.4] or [28, Definition 1.1]. A generating-cogenerating $n$-rigid subcategory $\mathcal{X}$ of ${\rm{{mod}}}(\Lambda)$ is called $n$-precluster tilting subcategory if it is functorially finite, $\tau_{n}(\mathcal{X})\subseteq\mathcal{X}$, and $\tau_{n}^{-}(\mathcal{X})\subseteq\mathcal{X}$ [31, Definition 3.2]. Moreover, if $\mathcal{X}$ admits an additive generator $M$, we say that $\mathcal{X}$ is a finite $n$-precluster tilting subcategory and $M$ is an $n$-precluster tilting module. By following Iyama-Solberg [31, Definition 3.4], Artin algebras which have $n$-precluster tilting modules are also called $\tau_{n}$-selfinjective.
@@
Now, by combining Iyama-Solberg correspondence and our results (Theorem 3.6), we can prove a result that gives us a new characterization of the class of Artin algebras having $n$-precluster tilting modules, i.e. $\tau_{n}$-selfinjective algebras. We need the following lemma which is of independent interest and follows from Lemmas 2.4 and 2.5.
@@
    The functor $\mathrm{Hom}_{\Lambda}(M,-):{\rm{{mod}}}(\Lambda)\to{\rm{{mod}}}(\Gamma)$ restricts to an equivalence $\mathrm{add}(M)\simeq\mathrm{proj}(\Gamma)$, which further restricts to another equivalence $\mathrm{inj}(\Lambda)\simeq\mathrm{proj}(\Gamma)\cap\mathrm{inj}(\Gamma)$, so $\mathrm{add}(\mathcal{Q})$ $\mathrm{proj}(\Gamma)\cap\mathrm{inj}(\Gamma)$.
@@
The statement $(a)$ can be proved as in [5, Lemma V.5.3]. For $(b)$, the fact that $M_{\Gamma}$ is projective follows from the duality $\mathrm{Hom}_{\Lambda}(-,M)$ between $\mathrm{add}(M)$ and $\mathrm{proj}(\Gamma^{op})$ and the fact that $_{\Lambda}\Lambda\in\mathrm{add}(M)$.
@@
For $(c)$, the two functors $\mathrm{Hom}_{\Gamma}(I,-)$ and $D\mathrm{Hom}_{\Gamma}(-,Q)$ from ${\rm{{mod}}}(\Gamma)$ to ${\rm{{mod}}}(\Lambda)$ are right exact (in fact exact), so by Watts’ theorem, one need to verify that they coincides with $M\otimes_{\Gamma}$- on the left regular module $_{\Lambda}\Lambda$, which is obvious by $Q=\mathrm{Hom}_{\Lambda}(M,D\Lambda)\simeq D(M)$ and $I\simeq\mathrm{Hom}_{\Gamma}(M,\Gamma)\simeq\nu^{-1}Q$. The existence of the adjoint triple follows from the usual tensor-Hom adjunction.
@@
By the first lemma of [3, Chapter III, Section 4, Page 48]. this equivalence can be extended to an equivalence ${\rm{{mod}}}(\Lambda)\simeq\mathrm{sub}^{2}(\mathcal{Q})$, as $Q$ is injective. This proves $(d)$.
For $(e)$, by [13, p. 17, Exercise 8] and [32, Theorem 3.2], the equivalence $\mathrm{inj}(\Lambda)\simeq\mathrm{add}(Q)$ can be extended to ${\mathrm{Inj}}(\Lambda)\simeq\mathrm{Add}(Q)$. Again the latter can be extended to ${\rm{Mod}}(\Lambda)\simeq\mathrm{Sub}^{2}(\mathcal{Q})$. ∎
@@
$(b)\Rightarrow(a)$. By Proposition 3.4, $\Gamma$ is an $n$-minimal Auslander-Gorenstein algebra and so by Iyama-Solberg correspondence [31, Theorem 4.5], there is an Artin algebra $\Lambda^{\prime}$ with $n$-precluster tilting $\Lambda^{\prime}$-module $M^{\prime}$ such that ${\rm{{mod}}}(\Lambda^{\prime})\simeq{\mathrm{Gproj}}^{\leq n-1}(\Gamma)\simeq{\rm{{mod}}}(\Lambda)$. Hence, $\Lambda$ and $\Lambda^{\prime}$ are also Morita equivalent.
@@
Lemmas 2.5 and 4.2, allowed us to generalize our results from finitely generated modules to all modules, see Theorems 3.6 and 4.3. Of course, we are not limited to these lemmas and can follow other ways.
@@
By [34, Proposition 3.5], there is a (locally Noetherian) Grothendieck category $\mathcal{A}$ and a functor $F:\mathcal{A}\longrightarrow{\rm{Mod}}(\Gamma)$ such that $F$ induces an equivalence between ${\mathrm{Inj}}(\mathcal{A})$ and ${\rm{Add}}(\mathcal{Q})$. By the proof of [34, Proposition 3.5], the functor $F$ is a composition of the right adjoint of a quotient functor with an exact functor and so $F$ is left exact. Hence, $F$ induces an equivalence between $\mathcal{A}$ and $\mathrm{Sub}^{2}(\mathcal{Q})$. On the other hand, by a similar argument like the proof of Proposition 3.1, ${\mathrm{GProj}}^{\leq n-1}(\Gamma)=\mathrm{Sub}^{2}(\mathcal{Q})$ and so ${\mathrm{GProj}}^{\leq n-1}(\Gamma)$ is an abelian category.